OpenStudy (anonymous):

find the equation of the perpendicular bisector of the line segment with end points (3,-7) (6,-2)

OpenStudy (anonymous):

hey remeber me?

OpenStudy (anonymous):

y+7 = -3/5 (x-3) 5(y+7) =-3 (x-3) 5y+35 = -3x+9 5y=-3x +9-35 y=(-3x-26)/5 y= -3/5 x -26/5 this is the equation of perpendicular bisector of line